OpenStudy (anonymous):

find the derivative f(x)=cotx/(1+cscx) iget -csc^2x-csc^3+cscxcot^2x/(1+cscx)^2 but my book says the answer is -cscx/(1+cscx) how ca get that?

OpenStudy (diamondboy):

use d quotient rule

OpenStudy (anonymous):

f'(x) = -sinx (1+sinx) - (cosx)(cosx) / (1+sinx)^2 = (-sinx - sin^2 x - cos^2 x) / (1+sinx)^2 = (-sinx -1) / (1+sinx)^2 = - (1+sinx) / (1+sinx)^2 = -1 / (1+sinx)

OpenStudy (anonymous):

I did useth qtientrule how ca isipliy

OpenStudy (diamondboy):

@misty1212 what do u think about this?

OpenStudy (diamondboy):

Hi are you still there?

OpenStudy (diamondboy):

I think I have figured it out

OpenStudy (diamondboy):

\[\cot^2(x)*\csc(x)-\csc^2(x)*\csc(x)+1/\csc^2(x)+1\]

OpenStudy (diamondboy):

\[(\csc(X)).\frac{ \cot^2(x)-\csc^2(x) }{ \csc(x)+1 }= \frac{ \csc(x) *-1 }{ \csc(x)+1 }\]

OpenStudy (diamondboy):

and dat is your answer

OpenStudy (anonymous):

Notice that the answer obtained by the posts above are the same \[ -\frac{1}{\sin (x)+1}=-\frac{1}{\frac{1}{\csc (x)}+1}=-\frac{1}{\frac{\csc (x)+1}{\csc (x)}}=-\frac{\csc (x)}{\csc (x)+1} \]